You need a to get a grade of C (70%) in this course. At this point in the course, there have been 1000 points and you have 65% of these 1000 points.
The only thing left in the course is the final, and you feel very confident
that you will do well. In fact, you think that you can get at least a 90% on
the final. How many points do you need the final to be worth to bring your
grade up to a C?

Found 2 solutions by Boreal, solver91311:
Answer by Boreal(15235)

650/1000 is the current situation
You are 50 points below a C now
On the final, if you get a 90, you are getting 90 points for every 100 awarded or .9x for every x awarded.
so your percentage will be 650+.9x/1000+x, and that equals 7/10
so 6500+9x=7000+7x
2x=500
x=250 points on the final, and you need 90% of them or 225
your score will be 875/1250=0.7

Answer by solver91311(24713)



You need 5% of 1000 points.


John

My calculator said it, I believe it, that settles it